# Learning Multi-Index Models with Neural Networks via Mean-Field Langevin Dynamics

### v1

@@ 4 Polynomial Time Convergence in the Riemannian Setting
In the above result, the constants in Assumption 4 can be identified as $\bar{\varepsilon}=0$ and $\bar{\Delta}=\operatorname{osc}(f)$ which is the oscillation of the log-density of $\mu^{*}$. Consequently, if the neurons in the teacher model are sufficiently present in all directions of the weight space, we get $\operatorname{osc}(f)=o(d)$; consider e.g. the extreme case $\mu^{*}=\tau$ which implies $f$ is constant. Interestingly, in the case of $k$-multi-index models, this condition implies that $k$ grows with dimension, ruling out the case $k=\mathcal{O}(1)$.
@@ A.1 Concentration Bounds
Definition 11.
[72] A random variable $x$ is $\sigma$-subGaussian if $\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{e^{\lambda(x-\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{x}}\right])}}}\right]\leq e^{\lambda^{2}\sigma^{2}/2}$ for all $\lambda\in\mathbb{R}$, and is $(\nu,\alpha)$-subexponential if $\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{e^{\lambda(x-\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{x}}\right])}}}\right]\leq e^{\lambda^{2}\nu^{2}/2}$ for all $\mathopen{}\mathclose{{}\left|{\lambda}}\right|\leq 1/\alpha$. If $x$ is $\sigma$-subGaussian, then
@@
Under Assumption 1, we have $\mathbb{E}_{S_{n}}\mathopen{}\mathclose{{}\left[y^{2}}\right]\lesssim 1$ with probability at least $1-2\exp(-\Omega(n))$.
Proof.  We have
@@ A.2 Approximating the Target Function
$\hat{\mathcal{J}}_{0}(\mu^{*})$ $=\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{\rho(\hat{y}(\boldsymbol{x};\mu^{*})-y)}}\right]$

### v2

@@ 4 Polynomial Time Convergence in the Riemannian Setting
In the above result, the constants in Assumption 4 can be identified as $\bar{\varepsilon}=0$ and $\bar{\Delta}=\operatorname{osc}(f)$ which is the oscillation of the log-density of $\mu^{*}$. Consequently, if the neurons in the teacher model are sufficiently present in all directions of the weight space, we get $\operatorname{osc}(f)=o(d)$; consider e.g. the extreme case $\mu^{*}=\tau$ which implies $f$ is constant. Interestingly, in the case of $k$-multi-index models, this condition implies that $k$ grows with dimension, ruling out the case $k=\mathcal{O}(1)$. We include a natural example of target functions of interest in the form of Proposition 8 in Appendix D.
@@ A.1 Concentration Bounds
Definition 11 ([72]).
A random variable $x$ is $\sigma$-subGaussian if $\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{e^{\lambda(x-\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{x}}\right])}}}\right]\leq e^{\lambda^{2}\sigma^{2}/2}$ for all $\lambda\in\mathbb{R}$, and is $(\nu,\alpha)$-subexponential if $\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{e^{\lambda(x-\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[{x}}\right])}}}\right]\leq e^{\lambda^{2}\nu^{2}/2}$ for all $\mathopen{}\mathclose{{}\left|{\lambda}}\right|\leq 1/\alpha$. If $x$ is $\sigma$-subGaussian, then
@@
Under Assumption 1, we have $\mathbb{E}_{S_{n}}\mathopen{}\mathclose{{}\left[y^{2}}\right]\lesssim 1$ with probability at least $1-n^{-q}$.
Proof.  By the local Lipschitzness of $g$, on the event of Lemma 12, we have
@@ A.2 Approximating the Target Function
$\hat{\mathcal{J}}_{0}(\mu^{*})$ $=\mathbb{E}_{S_{n}}\mathopen{}\mathclose{{}\left[\rho(\hat{y}(\boldsymbol{x};\mu^{*})-y)}\right]$
@@ C.2 Comparison with the Effective Dimension of [55]
Appendix D Examples of Target Functions in Proposition 8
In this section, we provide a natural example of a target function of the form given in Proposition 8. Suppose $\mathcal{W}=\mathbb{S}^{d-1}$, and $\Psi(\boldsymbol{x};\boldsymbol{w})=\mathopen{}\mathclose{{}\left\langle{\boldsymbol{w}},{\boldsymbol{x}}}\right\rangle^{2}$. Define $\mu^{*}=\mathcal{T}\#\tau$, where $\mathcal{T}(\boldsymbol{v})=\frac{\boldsymbol{A}^{1/2}\boldsymbol{v}}{\mathopen{}\mathclose{{}\left\|{\boldsymbol{A}^{1/2}\boldsymbol{v}}}\right\|}$, for some PSD matrix $\boldsymbol{A}$ to be defined later. Let
$y=\hat{y}(\boldsymbol{x};\mu^{*})=\mathopen{}\mathclose{{}\left\langle{\boldsymbol{x}},{\operatorname{\mathbb{E}}_{\boldsymbol{w}\sim\mu^{*}}\mathopen{}\mathclose{{}\left[\boldsymbol{w}\boldsymbol{w}^{\top}}\right]\boldsymbol{x}}}\right\rangle=\mathopen{}\mathclose{{}\left\langle{\boldsymbol{x}},{\mathbb{E}_{\boldsymbol{v}\sim\tau}\Big{[}\frac{\boldsymbol{A}^{1/2}\boldsymbol{v}\boldsymbol{v}^{\top}\boldsymbol{A}^{1/2}}{\|\boldsymbol{A}^{1/2}\boldsymbol{v}\|^{2}}\Big{]}\boldsymbol{x}}}\right\rangle.$
By defining $\boldsymbol{B}=d\cdot\mathbb{E}_{\boldsymbol{v}\sim\tau}\Big{[}\frac{\boldsymbol{A}^{1/2}\boldsymbol{v}\boldsymbol{v}^{\top}\boldsymbol{A}^{1/2}}{\|\boldsymbol{A}^{1/2}\boldsymbol{v}\|^{2}}\Big{]}$, we have $y=\frac{1}{d}\mathopen{}\mathclose{{}\left\|{\boldsymbol{B}^{1/2}\boldsymbol{x}}}\right\|^{2}$. Additionally, note that $\hat{y}(\boldsymbol{x};\tau)=\frac{1}{d}\mathopen{}\mathclose{{}\left\|{\boldsymbol{x}}}\right\|^{2}.$ Before proceeding further, we remark that the typical $\mathopen{}\mathclose{{}\left\|{\boldsymbol{x}}}\right\|$ should be of order $\sqrt{d}$ to get $\Theta(1)$ output, which is due to choosing the unit sphere as the weight space.
Next, we construct $\boldsymbol{A}$. Suppose distribution of $\boldsymbol{x}$ is such that with probability 1/2 we have $\boldsymbol{x}=\boldsymbol{e}_{1}$, the first standard basis vector. Let $\boldsymbol{A}=\mathop{\mathrm{missing}}{diag}(\lambda_{1},1,\ldots,1)$. Then,
$y=\hat{y}(\boldsymbol{e}_{1};\mu^{*})=\operatorname{\mathbb{E}}_{\boldsymbol{v}\sim\tau}\mathopen{}\mathclose{{}\left[\frac{\lambda_{1}dv_{1}^{2}}{1+(\lambda_{1}-1)v_{1}^{2}}}\right]\leq c\lambda_{1},$
where $c$ is an absolute constant. On the other hand, $\hat{y}(\boldsymbol{e}_{1};\tau)=1$. Choosing $\lambda_{1}=\frac{1}{d}$, we observe that for sufficiently large $d$, with probability at least 1/2 the distance $\mathopen{}\mathclose{{}\left|{\hat{y}(\boldsymbol{x};\mu^{*})-\hat{y}(\boldsymbol{x};\tau)}}\right|\geq C$ for some absolute constant $C>0$. This means that $\mu^{*}$ and $\tau$ are meaningfully different, and from an initialization of $\tau$ one needs to train the network (e.g. with MFLD) to recover $\mu^{*}$. It remains to find an estimate on $\mathcal{H}(\mu^{*}\,|\,\tau)$.
Let $\mathfrak{T}:\mathbb{R}^{d}\to\mathbb{S}^{d-1}$ denote the normalization mapping, i.e. $\mathfrak{T}(\boldsymbol{v})=\frac{\boldsymbol{v}}{\mathopen{}\mathclose{{}\left\|{\boldsymbol{v}}}\right\|}$. Then, note that $\tau=\mathfrak{T}\#\mathcal{N}(0,\mathbf{I}_{d})$, and $\mu^{*}=\mathfrak{T}\#\mathcal{N}(0,\boldsymbol{A})$. Therefore, by the data processing inequality,
$\mathcal{H}(\mu^{*}\,|\,\tau)\leq\mathcal{H}(\mathcal{N}(0,\boldsymbol{A})\,|\,\mathcal{N}(0,\mathbf{I}_{d}))=-\frac{d}{2}+\frac{\mathop{\mathrm{missing}}{tr}(\boldsymbol{A})}{2}-\frac{1}{2}\ln\det(\boldsymbol{A})\leq\frac{\ln d}{2},$
where $\lambda_{1}=\frac{1}{d}$. Therefore, $\bar{\Delta}=\frac{\ln d}{2}=o(d)$, and we can apply Theorem 10 to obtain polynomial convergence time, as desired.
Appendix E Numerical Simulation
In this section, we perform numerical simulations to verify the intuitions from Theorem 3. Specifically, we train a two-layer neural network with width $m=50$ and ReLU activation, where the first layer weights are initialized uniformly on the sphere, and fix the first half of the second layer coordinates at $+1/m$, and the second half at $-1/m$. The input follows the distribution $\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{\Sigma})$ (with an extra 1 appended for bias), where $\boldsymbol{\Sigma}=\mathop{\mathrm{missing}}{diag}(\boldsymbol{\sigma}^{2})$ with $\sigma_{1}^{2}=1$ and $\sigma_{i}^{2}=\frac{d_{\mathrm{eff}}-1}{d-1}$ for input dimension $d=50$. The labels are generated by a single-index model of the following form
$y=g(\mathopen{}\mathclose{{}\left\langle{\boldsymbol{e}_{1}},{\boldsymbol{x}}}\right\rangle)=\frac{\mathopen{}\mathclose{{}\left\langle{\boldsymbol{e}_{1}},{\boldsymbol{x}}}\right\rangle^{2}-1}{\sqrt{2}}.$
Therefore, the effective dimension from Definition 1 is exactly equal to $d_{\mathrm{eff}}$. We train the neural network using the squared loss with MFLA, with a stepsize of $0.1$, weight decay parameter $0.01$, temperature $0.001$.
Figure 2: Generalization gap measured by variying the effective dimension.
Figure 1(b) shows the test loss at the end of 200 iterations of MFLA for different numbers of training samples $n$ and effective dimension $d_{\mathrm{eff}}$. For each value of $n$ and $d_{\mathrm{eff}}$, we average the test loss over 5 independent runs with different realizations of data and initialization. In Figure 2 we measure the generalization gap, i.e. the average loss difference on the training set of $n$ samples, and a test set of $100000$ samples, at the end of 3000 iterations of training with MFLA. For this experiment, we try $n=100$, $n=200$, and $n=500$. As seen from both figures, $d_{\mathrm{eff}}$ controls the generalization gap and test loss, both of which decay with larger $n$.
The code to reproduce the experimental results is provided at: https://github.com/mousavih/MFLD-Learnability.